# Stochastic weight matrix dynamics during learning and Dyson Brownian motion

### v1

@@ Appendix A Dyson Brownian motion and the stochastic Coulomb gas
In this Appendix we summarise the main concepts of Dyson Brownian motion [2], see in particular the text book [8] (Ch. 7). For simplicity we use continuous-time notation, but when in doubt, expressions should be understood in the sense of $\rm\hat{I}$to calculus.
Consider a symmetric $N\times N$ matrix $X$ updated according to
$\frac{dX_{ij}}{dt}=K_{ij}(X)+\sqrt{A_{ij}}\eta_{ij},$ (A.1)
where $K_{ij}$ is the drift term, $A_{ij}$ encodes the strength of the noise, and $\eta_{ij}\sim{\cal N}(0,1)$. Its eigenvalues $x_{i}$ then evolve according to
@@
$\equiv K_{i}^{\rm(eff)}(x_{i})+\sqrt{2}g_{i}\eta_{i},$ (A.2)
where $K_{i}(x_{i})$ is the drift acting on the eigenvalues, the second term is the induced Coulomb term, and $\eta_{i}\sim{\cal N}(0,1)$. This result can be derived using second-order perturbation theory with a discrete time step $\delta t$, making the standard assumption that the drift and the noise correlator $\langle\eta_{ij}\eta_{kl}\rangle$ scale as $\delta t$ [8]. The noise strength satisfies $A_{ij}=A(1+\delta_{ij})$ and we denote the diagonal element as $\sqrt{A_{ii}}=\sqrt{2}g_{i}$. Strictly speaking the index on $g_{i}$ should be dropped, but we keep it for future convenience.
@@
$\partial_{t}P(\{x_{i}\},t)=\sum_{i=1}^{N}\partial_{x_{i}}\left[\left(g_{i}^{2}\partial_{x_{i}}-K_{i}^{\rm(eff)}(x_{i})\right)\right]P(\{x_{i}\},t).$ (A.3)
@@
$P_{s}(\{x_{i}\})=\frac{1}{Z}\prod_{i<j}\left|x_{i}-x_{j}\right|e^{-\sum_{i}V_{i}(x_{i})/g_{i}^{2}},$ (A.5)
@@
$V(x_{1},x_{2})=\frac{1}{2\sigma^{2}}\left[(x_{1}-\kappa)^{2}+(x_{2}-\kappa)^{2}\right].$ (A.8)
@@
$P(S)=\frac{S}{2\sigma^{2}}e^{-S^{2}/(4\sigma^{2})}.$ (A.11)
@@
$\rho(x)=\left\langle\frac{1}{N}\sum_{i=1}^{N}\delta(x-x_{i})\right\rangle.$ (A.14)
@@
$\rho(x)=\frac{e^{-x^{2}/(2\sigma^{2})}}{4\sqrt{\pi}\sigma}\left[2e^{-x^{2}/(2\sigma^{2})}+\sqrt{2\pi}\frac{x}{\sigma}\mbox{Erf}\left(\frac{x}{\sqrt{2}\sigma}\right)\right].$ (A.15)
@@ Appendix B Weight matrix initialisation
Consider the $M\times N$ random matrix $W$ at initialisation, with elements $W\sim{\cal N}(0,\sigma_{\rm in}^{2})$. Without loss of generality we take $N\leq M$; if this is not the case, simply exchange $W$ and $W^{T}$. The symmetric $N\times N$ matrix $X=W^{T}W$ has $N$ nonzero eigenvalues $x_{i}=\xi_{i}^{2}$ ($i=1,\ldots,N$), where $\xi_{i}$ are the singular values of $W$. The distribution of eigenvalues $x$ of $X$ is given by the Marchenko-Pastur distribution,
$P_{\rm MP}(x)=\frac{1}{2\pi\sigma_{\rm in}^{2}Mrx}\sqrt{(x_{+}-x)(x-x_{-})},$ (B.1)
@@ Appendix C Non-universal part of the linear scaling relation
In Sec. 2, we demonstrated that the Coulomb gas potential $V_{i}(x_{i})/g_{i}^{2}$ is the product of a universal and a non-universal factor, see Eqs. (2.16, 2.18). Here we consider the non-universal factor in the case of the restricted Boltzmann machine and sketch a derivation of its parameter dependence.
We are interested in the variance of the gradient of the loss function, ${\rm Var}\left(\delta W\right)$ or ${\rm Var}\left(\delta X\right)$. We write the gradient in the form, see Eq. (3.5) and putting $\sigma_{h}^{2}=1$,
$\frac{\delta{\cal L}}{\delta W}W^{T}=\left(C^{\rm target}-C^{\rm RBM}\right)(WW^{T}).$ (C.1)
@@
$\left\langle C^{\rm RBM}\right\rangle=K^{-1},\qquad\qquad K=\mu^{2}1\!\!1-WW^{T}.$ (C.4)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)=\left(\left\langle C^{\rm RBM}C^{\rm RBM}\right\rangle-\left\langle C^{\rm RBM}\right\rangle^{2}\right)(WW^{T})(WW^{T}).$ (C.6)
In the Gaussian RBM only the second moment is nontrivial, see Eq. (C.4). Hence
$\left\langle C^{\rm RBM}C^{\rm RBM}\right\rangle-\left\langle C^{\rm RBM}\right\rangle^{2}=2K^{-1}K^{-1}.$ (C.7)
@@
$K^{-1}\to\frac{1}{\mu^{2}-x},\qquad\qquad WW^{T}\to x,$ (C.8)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)\sim\frac{2x^{2}}{(\mu^{2}-x)^{2}}.$ (C.9)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)\sim\frac{2\left(\mu^{2}-\kappa\right)^{2}}{\kappa^{2}}.$ (C.10)
@@
$\tilde{g}_{i}^{2}\sim\kappa_{i}^{2}\Omega_{i}^{2},$ (C.11)
or, see Eq. (2.18),
$\sigma_{i}^{2}=\frac{\alpha}{|{\cal B}|}\kappa_{i}^{2}\Omega_{i}.$ (C.12)
@@
$x_{n+1}=x_{n}-\alpha\omega x_{n}+\sqrt{2\alpha\gamma}\eta_{n},\qquad\qquad\gamma=\frac{\alpha A}{|{\cal B}|},\qquad\qquad\langle\eta_{n}\eta_{m}\rangle=\delta_{nm}.$ (C.13)
It is solved by (taking $x_{0}=0$)
$x_{n}=\sqrt{2\alpha\gamma}\sum_{i=0}^{n-1}(1-\alpha\omega)^{n-1-i}\eta_{i}.$ (C.14)
@@
$\lim_{n\to\infty}\langle x_{n}^{2}\rangle=\frac{\alpha A}{|{\cal B}|}\frac{1}{\omega}\frac{1}{1-\alpha\omega/2}.$ (C.15)

### v2

@@ Appendix A Dyson Brownian motion and the stochastic Coulomb gas
In this Appendix we summarise the main concepts of Dyson Brownian motion Dyson-4 , see in particular the text book Meh2004 (Ch. 7). For simplicity we use continuous-time notation, but when in doubt, expressions should be understood in the sense of Itô calculus.
Consider a symmetric $N\times N$ matrix $X$, whose matrix elements are updated according to
$\frac{dX_{ij}}{dt}=K_{ij}(X)+\sqrt{A_{ij}}\eta_{ij},$ (62)
where $K_{ij}$ is the drift term, $A_{ij}$ encodes the strength of the noise, and $\eta_{ij}\sim{\cal N}(0,1)$, all of which are symmetric. Its eigenvalues $x_{i}$ then evolve according to
@@
$\equiv K_{i}^{\rm(eff)}(x_{i})+\sqrt{2}g_{i}\eta_{i},$ (63)
where $K_{i}(x_{i})$ is the drift acting on the eigenvalues, the second term is the induced Coulomb term, and $\eta_{i}\sim{\cal N}(0,1)$. This result can be derived using second-order perturbation theory with a discrete time step $\delta t$, making the standard assumption that the drift and the noise correlator $\langle\eta_{ij}\eta_{kl}\rangle$ scale as $\delta t$ Meh2004 . The noise strength satisfies $A_{ij}=A(1+\delta_{ij})$ and we denote the diagonal element as $\sqrt{A_{ii}}=\sqrt{2}g_{i}$. Strictly speaking the index on $g_{i}$ should be dropped, but we keep it for future convenience. A derivation in the context of the teacher-student model is given in App. D.
@@
$\partial_{t}P(\{x_{i}\},t)=\sum_{i=1}^{N}\partial_{x_{i}}\left[\left(g_{i}^{2}\partial_{x_{i}}-K_{i}^{\rm(eff)}(x_{i})\right)\right]P(\{x_{i}\},t).$ (64)
@@
$P_{s}(\{x_{i}\})=\frac{1}{Z}\prod_{i<j}\left|x_{i}-x_{j}\right|e^{-\sum_{i}V_{i}(x_{i})/g_{i}^{2}},$ (66)
@@
The presentation above assumes that the potential in the Coulomb gas is separable, see Eqs. (65, 66). If this is not possible, e.g. due to eigenvalues mixing non-linearly in the drift terms, a more general potential appears, as
$\sum_{i}V_{i}(x_{i})/g_{i}^{2}\to\overline{V}(x_{1},x_{2},\ldots,x_{N}),$ (68)
with
$K_{i}=-g_{i}^{2}\frac{\partial\overline{V}(x_{1},x_{2},\ldots,x_{N})}{\partial x_{i}}.$ (69)
While this will complicate the analysis and lead to a richer loss function landscape, the implications are expected to remain.
@@
$V(x_{1},x_{2})=\frac{1}{2\sigma^{2}}\left[(x_{1}-\kappa)^{2}+(x_{2}-\kappa)^{2}\right].$ (71)
@@
$P(S)=\frac{S}{2\sigma^{2}}e^{-S^{2}/(4\sigma^{2})}.$ (74)
@@
$\rho(x)=\left\langle\frac{1}{N}\sum_{i=1}^{N}\delta(x-x_{i})\right\rangle.$ (77)
@@
$\rho(x)=\frac{e^{-x^{2}/(2\sigma^{2})}}{4\sqrt{\pi}\sigma}\left[2e^{-x^{2}/(2\sigma^{2})}+\sqrt{2\pi}\frac{x}{\sigma}\mbox{Erf}\left(\frac{x}{\sqrt{2}\sigma}\right)\right].$ (78)
@@ Appendix B Weight matrix initialisation
Consider the $M\times N$ random matrix $W$ at initialisation, with elements $W\sim{\cal N}(0,\sigma_{\rm in}^{2})$. Without loss of generality we take $N\leq M$; if this is not the case, simply exchange $W$ and $W^{T}$. The symmetric $N\times N$ matrix $X=W^{T}W$ has $N$ non-zero eigenvalues $x_{i}=\xi_{i}^{2}$ ($i=1,\ldots,N$), where $\xi_{i}$ are the singular values of $W$. The distribution of eigenvalues $x$ of $X$ is given by the Marchenko-Pastur distribution,
$P_{\rm MP}(x)=\frac{1}{2\pi\sigma_{\rm in}^{2}Mrx}\sqrt{(x_{+}-x)(x-x_{-})},$ (79)
@@ Appendix C Non-universal part of the linear scaling relation
In Sec. II, we demonstrated that the Coulomb gas potential $V_{i}(x_{i})/g_{i}^{2}$ is the product of a universal and a non-universal factor, see Eqs. (16, 18). Here we consider the non-universal factor in the case of the restricted Boltzmann machine and sketch a derivation of its parameter dependence.
We are interested in the variance of the gradient of the loss function, ${\rm Var}\left(\delta W\right)$ or ${\rm Var}\left(\delta X\right)$. We write the gradient in the form, see Eq. (24) and putting $\sigma_{h}^{2}=1$,
$\frac{\delta{\cal L}}{\delta W}W^{T}=\left(C^{\rm target}-C^{\rm RBM}\right)(WW^{T}).$ (81)
@@
$\left\langle C^{\rm RBM}\right\rangle=K^{-1},\qquad\qquad K=\mu^{2}1\!\!1-WW^{T}.$ (84)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)=$
$\left(\left\langle C^{\rm RBM}C^{\rm RBM}\right\rangle-\left\langle C^{\rm RBM}\right\rangle^{2}\right)(WW^{T})(WW^{T}).$ (86)
In the Gaussian RBM only the second moment is non-trivial, see Eq. (84). Hence
$\left\langle C^{\rm RBM}C^{\rm RBM}\right\rangle-\left\langle C^{\rm RBM}\right\rangle^{2}=2K^{-1}K^{-1}.$ (87)
@@
$K^{-1}\to\frac{1}{\mu^{2}-x},\qquad\qquad WW^{T}\to x,$ (88)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)\sim\frac{2x^{2}}{(\mu^{2}-x)^{2}}.$ (89)
@@
${\rm Var}\left(\frac{\delta{\cal L}}{\delta W}W^{T}\right)\sim\frac{2\left(\mu^{2}-\kappa\right)^{2}}{\kappa^{2}}.$ (90)
@@
$\tilde{g}_{i}^{2}\sim\kappa_{i}^{2}\Omega_{i}^{2},$ (91)
or, see Eq. (18),
$\sigma_{i}^{2}=\frac{\alpha}{|{\cal B}|}\kappa_{i}^{2}\Omega_{i}.$ (92)
@@
$x_{n+1}=x_{n}-\alpha\omega x_{n}+\sqrt{2\alpha\gamma}\eta_{n},$ (93)
where $\gamma=\alpha A/|{\cal B}|$ and $\langle\eta_{n}\eta_{m}\rangle=\delta_{nm}$. It is solved by (taking $x_{0}=0$)
$x_{n}=\sqrt{2\alpha\gamma}\sum_{i=0}^{n-1}(1-\alpha\omega)^{n-1-i}\eta_{i}.$ (94)
@@
$\lim_{n\to\infty}\langle x_{n}^{2}\rangle=\frac{\alpha A}{|{\cal B}|}\frac{1}{\omega}\frac{1}{1-\alpha\omega/2}.$ (95)
@@
Appendix D Derivation of the eigenvalue equation
In this appendix we add a brief derivation of the eigenvalue equation including the Coulomb term, starting with the matrix update, in the context of the teacher-student model. This follows closely text book derivations Meh2004 , but may be useful to those not familiar with this topic.
We consider a $2\times 2$ symmetric matrix $W$, updated according to
$W_{ij}\to W_{ij}^{\prime}=W_{ij}-\alpha\frac{\delta{\cal L}}{\delta W_{ij}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{A_{ij}}\eta_{ij},$ (96)
where $\alpha$ is the learning rate and $|{\cal B}|$ the batch size. Both the noise $\eta_{ij}\sim{\cal N}(0,1)$ and strength $A_{ij}$ are symmetric matrices. The eigenvalues of $W$ are denoted as $\lambda_{1,2}$ and the teacher values are $\kappa_{1,2}$.
We assume that at the current step $W$ is diagonal (or has been diagonalised) and consider one update. The drift is then $-\alpha(\lambda_{1,2}-\kappa_{1,2})$ and the updated matrix reads
$W^{\prime}=\begin{pmatrix}\lambda_{1}-\alpha(\lambda_{1}-\kappa_{1})+(\alpha\sqrt{|{\cal B}|})\sqrt{A_{11}}\eta_{11}&(\alpha/\sqrt{|{\cal B}|})\sqrt{A_{12}}\eta_{12}\\ (\alpha/\sqrt{|{\cal B}|})\sqrt{A_{12}}\eta_{12}&\lambda_{2}-\alpha(\lambda_{2}-\kappa_{2})+(\alpha/\sqrt{|{\cal B}|})\sqrt{A_{22}}\eta_{22}\end{pmatrix}.$ (97)
To compute the eigenvalues of $W^{\prime}$, we follow the standard power counting rules and treat the drift as ${\cal O}(\epsilon)$ and the noise terms as ${\cal O}(\sqrt{\epsilon})$. Expanding to ${\cal O}(\epsilon)$ then yields the eigenvalues of $W^{\prime}$,
$\lambda_{1}^{\prime}=\lambda_{1}-\alpha(\lambda_{1}-\kappa_{1})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A_{12}\eta_{12}^{2}}{\lambda_{1}-\lambda_{2}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{A_{11}}\eta_{11},$ (98)
$\lambda_{2}^{\prime}=\lambda_{2}-\alpha(\lambda_{2}-\kappa_{2})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A_{12}\eta_{12}^{2}}{\lambda_{2}-\lambda_{1}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{A_{22}}\eta_{22}.$ (99)
Taking a noise average in the Coulomb term, i.e., replacing $\eta_{12}^{2}$ by $1$, denoting $\eta_{11}=\eta_{1},\eta_{22}=\eta_{2}$, and using that $A_{ij}=A(1+\delta_{ij})$ (or $A_{11}=A_{22}=2A,A_{12}=A_{21}=A$Meh2004 , then yields
$\lambda_{1}^{\prime}=\lambda_{1}-\alpha(\lambda_{1}-\kappa_{1})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{1}-\lambda_{2}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{1},$ (100)
$\lambda_{2}^{\prime}=\lambda_{2}-\alpha(\lambda_{2}-\kappa_{2})+\frac{\alpha^{2}}{|{\cal B}|}\frac{A}{\lambda_{2}-\lambda_{1}}+\frac{\alpha}{\sqrt{|{\cal B}|}}\sqrt{2A}\eta_{2}.$ (101)
This is the standard equation for the eigenvalues, including the Coulomb and stochastic terms. As mentioned, a detailed discussed of the various steps and an extension to $N\times N$ matrices can be found in the text book Meh2004 .